Bruce Gerlinger is a pickpocket who has spent 3/5 of his life in jail. He is there currently, with 2 more years to serve. If at the end of those 2 years he stays out of jail entirely for as many years as he will have spent in jail, he will be 44 years old. How old is he now?
Answer by CharlesG2(834)

Bruce Gerlinger is a pickpocket who has spent 3/5 of his life in jail. He is there currently, with 2 more years to serve. If at the end of those 2 years he stays out of jail entirely for as many years as he will have spent in jail, he will be 44 years old. How old is he now?
let x = age now
(3/5)x + 2 = time in jail
(3/5)x + 2 = time out of jail
(3/5)x + 2 + (3/5)x + 2 = 44
(6/5)x + 4 = 44
(6/5)x = 40
x = 5/6 * 40
x = 200/6 = 33 1/3 to nearest year he is 33 years old now
check:
(3/5)(200/6) + 2 = 600/30 + 2 = 60/3 + 2 = 20 + 2 = 22
2 * 22 = 44